A gas of 4.68 moles of monatomic molecules is held at a constant temperature of 47.7C as it expands. As it expands, 29.0 Joules of heat enter the gas. What is the change in entropy of the gas?
Answer #1

dS = dQ/T = 29 / (273+47.7) = 0.0904 J/K per 4.68moles => 19.32 mJ /K-mol